A 0.3146-g sample of a mixture of NaCl(s) and KBr(s) was dissolved in water. The resulting solution required 53.60 mL of 0.08765 M AgNO3(aq) to precipitate the Cl
Answer #1
Mol(Ag+) = molarity * volume
= (0.08765mol/1000ml)*53.6ml
= 4.69804x10-3 mol
These moles are equal to the total number of mols of NaCl and KBr.
Let the percentage by mass for:
NaCl be X%
Therefore the % by mass for KBr becomes (100-X)%
Mol(NaCl) = mass of NaCl/Formula weight of NaCl
= (0.3146*X/100)/58.44
= 5.3833*10--5X
Mol(KBr) = mass of KBr/Formula weight of KBr
= (0.3146*(100-X)/100)/119.01
= 2.6435*10-5(100-X)

Since mol(NaCl) + mol(KBr) = mol(Ag+)
5.3833*10--5X + 2.6435*10-5(100-X) = 4.69804x10-3
2.7398*10--5X + 2.6435*10-3 = 4.69804x10-3
2.7398*10--5X = 2.05454*10-3
X = 74.99%
Therefore, the sample contains 74.99% and 25.01% of NaCl and KBr respectively